OpenStudy (kaley246):

Graph the two equations to solve the system. What is the ordered pair solution? x + y = 6 and x - y = 4 (Points : 1) (4, 2) (6, 2) (5, 1) no solution

OpenStudy (kaley246):

@AnswerMyQuestions

OpenStudy (anonymous):

Use this graphing calculator: https://www.desmos.com/calculator

OpenStudy (anonymous):

Plug in the lines, and you get your solution. :)

OpenStudy (kaley246):

it just make a x

OpenStudy (anonymous):

Yes, but what is the ordered pair where the lines touch?

OpenStudy (kaley246):

like the wnd of the lines

OpenStudy (kaley246):

@AnswerMyQuestions

OpenStudy (anonymous):

?

OpenStudy (anonymous):

https://www.desmos.com/calculator/p1fzft6oiu

OpenStudy (anonymous):

What is the point that they intersect at?

OpenStudy (kaley246):

there is a person that keeps on saying bad words

OpenStudy (kaley246):

i dont know

OpenStudy (kaley246):

what the points are its 0

OpenStudy (kaley246):

@robtobey

OpenStudy (anonymous):

WHERE DO THE LINES TOUCH?

OpenStudy (kaley246):

idont know can u do it

OpenStudy (anonymous):

smh

OpenStudy (kaley246):

what does that mean

OpenStudy (kaley246):

0

OpenStudy (kaley246):

5,1

OpenStudy (kaley246):

@AnswerMyQuestions

OpenStudy (kaley246):

@pielover123

OpenStudy (anonymous):

Your right

OpenStudy (kaley246):

5,1 that right

OpenStudy (kaley246):

@pielover123

OpenStudy (kaley246):

Graph the two equations to solve the system. What is the solution? 1/2 x - 2y = 4 and 4y - x = -8 (Points : 1) (-2, 0) no solution (3, 4) infinitely many solutions

OpenStudy (anonymous):

B. No solution

OpenStudy (kaley246):

how is it b

OpenStudy (kaley246):

@pielover123

OpenStudy (kaley246):

Graph the two equations to solve the system. What is the solution? y = 2x - 4 and 2x + 3y = -12 (Points : 1) (0, -4) (4, 4) no solution infinitely many solutions

OpenStudy (anonymous):

~_~ Im not 100% sure but im going with infinitely many solutions

OpenStudy (kaley246):

Graph the two equations to solve the system. What is the solution? x + y = -8 and x - y = 8 (Points : 1) (0, -8) (-12, 4) (-4, -4) no solution

OpenStudy (kaley246):

@pielover123

OpenStudy (kaley246):

@pielover123

OpenStudy (anonymous):

(-4, -4)

OpenStudy (kaley246):

Graph the two equations to solve the system. What is the solution? 3x - 2y = 6 and 5x + 3y = -9 (Points : 1) (4, 3) (5, 4.5) (0, -3) infinitely many solutions

OpenStudy (anonymous):

infinitely many solutions

OpenStudy (kaley246):

Is the system consistent or inconsistent? 2x + y = 7 and x - 2y = 7 (Points : 1) consistent inconsistent

OpenStudy (kaley246):

@pielover123

OpenStudy (kaley246):

@pielover123

OpenStudy (kaley246):

@iGreen

OpenStudy (kaley246):

@SaltLife23

OpenStudy (kaley246):

@phi

OpenStudy (kaley246):

@mayataylor

OpenStudy (kaley246):

@LittleRedRiding

OpenStudy (kaley246):

@Ashes.boom

OpenStudy (kaley246):

@asif_cheeni

OpenStudy (kaley246):

@CheesecakeKitten

OpenStudy (kaley246):

@gibbs

OpenStudy (kaley246):

@sidney_girl1419

OpenStudy (kaley246):

@pooja195

OpenStudy (gibbs):

idk

OpenStudy (kaley246):

@sleepyjess

OpenStudy (kaley246):

@cookiimonster27

OpenStudy (kaley246):

@AnswerMyQuestions

OpenStudy (kaley246):

@welshfella

OpenStudy (kaley246):

@phi

OpenStudy (kaley246):

@sleepyjess @slinley

OpenStudy (kaley246):

@andrea238

OpenStudy (kaley246):

Is the system consistent or inconsistent? 2x + y = 7 and x - 2y = 7 (Points : 1) consistent inconsistent

OpenStudy (kaley246):

@ccokiens